OpenStudy (anonymous):

A = set of all English words. R = {(x,y)∈AxA | the word y occurs at least as late in alphabetical order as the word x}. What does this even mean???

OpenStudy (anonymous):

You should think of R as a type of ordering relation, like \(<,\leq,>\geq \)

OpenStudy (anonymous):

In a way it is basically \(x \geq y\), where every word is sorted in alphabetical order, and \(x\) and \(y\) are some natural number which represent some words ordinal value.

OpenStudy (anonymous):

what does *occur at least as late * mean?

OpenStudy (anonymous):

ok how about let (x,y) = (cat, batman). Is it in R?

OpenStudy (usukidoll):

I see a Cartesian product

OpenStudy (anonymous):

Yes, it's a Cartesian product with a condition whose meaning I don't quite understand

OpenStudy (anonymous):

at least as late means it is the same word or it comes after the word in alphabetical order.

OpenStudy (anonymous):

\( A\times A \) is the set of any two words, such at \((\text{dog},\text{cat})\).

OpenStudy (anonymous):

\( (\text{dog},\text{cat})\notin R \) but \( (\text{cat},\text{dog})\in R \) also \( (\text{cat},\text{cat})\in R \)

OpenStudy (anonymous):

The second word must be the same or comes after the first word, so shouldn't (cat,dog) be in R instead?

OpenStudy (anonymous):

What? \((\text{dog},\text{cat})\in A\times A\)

OpenStudy (anonymous):

oh wait nvm. I just repeated what you said earlier.

OpenStudy (anonymous):

So (cat, batman) isn't in R, but (batman, cat) is right?

OpenStudy (usukidoll):

\[A \times B =[(a,b) : a \in A \land b \in B]\] there are elements in A and B it's an ordered pair...

OpenStudy (usukidoll):

so I see two A's won't that be elements in A and elements in A again?!

OpenStudy (anonymous):

That's correct.

OpenStudy (usukidoll):

so if A = set of all english words then we have the set of all english words and the set of all english words... like mouse mouse in A

OpenStudy (anonymous):

the components need not to be the same. (mouse, cat) is certainly in AxA

OpenStudy (usukidoll):

hmm late for y ... like the ending of the alphabet worm xray yarn zebra

OpenStudy (usukidoll):

(digimon, pokemon)

OpenStudy (anonymous):

(UsukiDoll, Wio) :DD

OpenStudy (anonymous):

anyway, thanks for the explanation @Wio